A closed-end fund’s NAV is $71 at the beginning of the year and $31 at the end of the year. At the beginning of the year the fund was selling at a 2% premium to NAV, at the end of the year it was selling at a 2% discount to NAV. The fund paid year-end distributions of income and capital gains of $1. What is the rate of the return to an investor, who bought the fund at the beginning of the year?

EXCEL PLEASE
Answer #1
Particulars Sl No Amount($)
beginning NAV a 71
premium to NAV (%) b 2%
purchase price to an investor c=a*(1+b) 72.42
year end distributions d 1
closing NAV e 31
disount to NAV (%) f 2%
sale price to the investor g=e*(1-f) 30.38
rate of return to the investor (%) h=(g+d-c)/c -56.67%